It takes 2.54ms for the current in an LR circuit to increase from zero to 0.80 its maximum value.

Determine the time constant of the circuit.

Determine the resistance of the circuit if 30.0 mH.
Answer #1

Apply I = Imax(1 - e-t/t')

.8 = (1 - e-(2.54 X 10-3)/t')

.2 = e-(2.54 X 10-3)/t'

-1.609 = -(2.54 X 10-3)/t'

t' = 1.58 X 10-3 sec

t' = L/R

(1.58 X 10-3) = (30 X 10-3)/R

R = 19.0 Ohms